OpenStudy (anonymous):

A worker does 500J of wark on a 10kg box. If the box transfers 375J of heat to the floor through the friction between the box and the floor, what is the velocity of the box after the work has been done on it?

OpenStudy (anonymous):

\[(500-300)=(1/2)(Mass)(Velocity)^{2}\]

OpenStudy (anonymous):

should be 500-375 on the LHS but the rest is good. Just one tiny mistake in rewritting the info from the question.

OpenStudy (anonymous):

work done=Kinetic energy+Heat energy \[500=0.5(10)v^2+375\]\[v=\pm5m/s\]